# PUMA: Efficient Continual Graph Learning with Graph Condensation

### v1

@@ Abstract
When handling streaming graphs, existing graph representation learning models encounter a catastrophic forgetting problem, where previously learned knowledge of these models is easily overwritten when learning with newly incoming graphs. In response, Continual Graph Learning (CGL) emerges as a novel paradigm enabling graph representation learning from static to streaming graphs. Our prior work, Condense and Train (CaT) [1] is a replay-based CGL framework with a balanced continual learning procedure, which designs a small yet effective memory bank for replaying data by condensing incoming graphs. Although the CaT alleviates the catastrophic forgetting problem, there exist three issues: (1) The graph condensation algorithm derived in CaT only focuses on labelled nodes while neglecting abundant information carried by unlabelled nodes; (2) The continual training scheme of the CaT overemphasises on the previously learned knowledge, limiting the model capacity to learn from newly added memories; (3) Both the condensation process and replaying process of the CaT are time-consuming. In this paper, we propose a PsUdo-label guided Memory bAnk (PUMA) CGL framework, extending from the CaT to enhance its efficiency and effectiveness by overcoming the above-mentioned weaknesses and limits. To fully exploit the information in a graph, PUMA expands the coverage of nodes during graph condensation with both labelled and unlabelled nodes. Furthermore, a training-from-scratch strategy is proposed to upgrade the previous continual learning scheme for a balanced training between the historical and the new graphs. Besides, PUMA uses a one-time prorogation and wide graph encoders to accelerate the graph condensation and the graph encoding process in the training stage to improve the efficiency of the whole framework. Extensive experiments on four datasets demonstrate the state-of-the-art performance and efficiency over existing methods. The code has been released in https://github.com/superallen13/PUMA.
@@ 1 Introduction
Graph representation learning aims to model the graph-structured data, typically in recommender systems [2], traffic prediction [3] and protein function prediction [4]. Generally, graphs are treated as static data in traditional graph representation learning, where a model is fixed once it has been trained. However, for many scenarios in the real world (e.g., social networks, cation networks and knowledge graphs), graph data are changing and evolving in a streaming manner [5, 6, 7]. Recently, the Continual graph learning (CGL) has emerged to handle the streaming graph by adapting the static graph neural networks (GNNs). In CGL, the most significant challenge is how to address a catastrophic forgetting problem, where a model easily forgets the previously learned knowledge while overemphasising on the incoming graphs [8, 9, 10, 1].
@@
Although our previous CaT framework [1] has demonstrated that the incoming graphs can be successfully condensed to an extremely small scale, for example, 0.1% of the original size, for the continual learning, it remains a few challenges in further improving the efficacy and efficiency. (1) Neglected unlabelled nodes: In the previous condensation process, nodes without label will be considered as the target for condensation rather than just supporting the message passing for the labelled nodes, potentially overlooking the valuable information from unlabelled nodes. (2) Imbalanced learned knowledge: Since previously generated memories will be repeatedly replayed during the continual learning, historical knowledge is easier to retain while the perception of new information in incoming graphs is hindered. Figure 1 shows this situation for CGL that the CaT model converges faster and better for the knowledge from previous Task 0 with a smaller loss than the newly incoming Task 1. This is due to the fact that neural networks are prone to learn shortcuts from training data [12]. (3) Low efficiency: Both the condensation and the CGL training are inefficient, because of repetitive computations, slow convergence speed of condensation and unnecessary message passing in the GNN training. The condensation process requires to repeatedly map the incoming graphs to different embedding spaces, which causes a large number of recalculations for message passing among node neighbourhoods. Besides, condensed graph converging is impeded by narrow graph encoders (e.g., encoders with 512-dimensional hidden layers). Furthermore, message aggregations of neighbours on the edge-free replayed memories is redundant, which hinders the efficiency of CGL model training.
@@
Figure 1: Loss and accuracy comparisons for the first two tasks when training replayed graph of Task 1 while replaying the replayed graph of Task 0 at the first 50 steps after the replayed graph of Task 0 has been learned 500 steps. The continual (dotted lines) and retrained (solid lines) loss values for Task 0 (red) and Task 1 (blue) are shown, providing a visual representation of an optimisation problem.
@@
In light of the above challenges in the CaT method for CGL, this paper proposes a novel PseUdo-label guided Memory bAnk (PUMA) CGL framework to extend the CaT in a more effective and efficient manner. Firstly, the PUMA framework develops a pseudo-labelling to integrate data from unlabelled nodes, enhancing the informativeness of the memory bank, which addresses the issue of neglected unlabelled nodes. Secondly, to tackle the inflexible historical knowledge problem, a retraining strategy is devised. This involves initialising the entire model before replaying to balance the learned knowledge among different tasks for more effective decision boundaries as well as ensuring a more stable learning process. Figure 1 represents the optimisation process is smother, and the converge speed is quicker after using retraining scheme. Lastly, the repetitive message passing calculation in the condensation process is simplified via a one-time propagation method, which aggregates messages from neighbouring nodes across the entire incoming graph one time, and it can be stored for reuse, significantly reducing the computational effort. Furthermore, wide graph encoders including more neurons are developed to accelerate convergence during condensation, enhancing the update efficiency of the condensed memories. Finally, to improve the training efficiency of the CGL model, the training of a multilayer perceptron (MLP) in lieu of message-passing GNNs is proposed. Due to its edge-free nature, PUMA enables the MLP to learn feature extractions and utilises a GNN to inference on the graphs with edges. These solutions collectively enhance the efficacy and efficiency of the graph condensation-based CGL framework.
@@ 2.1 Graph Neural Networks
Graph Neural Networks (GNNs) are effective tools for graph-based tasks [13, 14, 15, 16]. GCN [17] employs Laplacian normalisation for message propagation. GAT [18] is proposed to use the attention mechanism [19] for message passing. SGC [20] simplifies GCN by removing the non-linear activation operation. GraphSAGE [21] uses node sampling to deal with large-scale graph representation learning.
@@ 2.2 Continual Graph Learning
Continual graph learning is a task for handling streaming graph data. Continual learning has been studied in computer vision [22, 23, 24, 25]. In graph area, CGL methods can be categorised into three branches: regularisation [9], replay- [8, 10], and architecture-based [11] methods. TWP [9] preserves the topological information of historical graphs by adding regularisation. HPNs [11] redesigns the GNN architecture to 3-layer prototypes for representation learning. ER-GNN [8] integrates memory-replay by storing representative nodes. Sparsified subgraph memory (SSM) [10] stores the sparisified subgraphs by dropping edges and nodes randomly or basing on the node degree. Differing from SSM, which sparsifies the incoming graph based on the degree or randomly selected central nodes’ neighbours, SEM-curvature [26] employs balanced Forman curvature to compute edges containing the most informative content to maximise structural information within the subgraph. Considering the inter connections between tasks, the structure shift also leads to catastrophic forgetting problem. Structural shift risk mitigation (SSRM) [27] uses regularisation-based techniques to mitigate this issue. Two recent benchmarks [28, 29] have been developed for CGL.
@@ 2.3 Graph Condensation
Dataset condensation generates a small and synthetic dataset to replace the original dataset and to train a model with similar performance. Dataset condensation has been applied in computer vision [30, 31, 32, 33, 34, 35]. Recently, gradient matching has been applied to graph condensation, such as GCond [36], DosCond [37] and MCond [38]. DM [34] aims to learn synthetic samples which have similar distribution with the original dataset to mimic sampling methods [39, 40, 41]. GCDM [42] uses the distribution matching for graph condensation. SFGC [43] uses a training trajectory meta-matching scheme to condensed the original graph into edge-free graphs. SGDD [44] broadcasts the structure information of original graph to condensed graph in the spectral graph theory perspective.
Recent attempts on computer vision have directly applied dataset condensation to continual learning [45, 46, 47, 48, 49]. Our previous work CaT [1] applies graph condensation to graph continual learning while ensuring the balanced training.
@@ 3 Preliminary
In the following, a bold lowercase letter denotes a vector, a bold uppercase letter denotes a matrix, a general letter denotes a scalar, and a scripted uppercase letter denotes a set.
3.1 Graph
For a node classification problem, a graph is denoted as $\mathcal{G}=\{\boldsymbol{A},\boldsymbol{X},\boldsymbol{Y}\}$, where $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ is the $d$-dimensional feature matrix for $n$ nodes, and the adjacency matrix $\boldsymbol{A}\in\mathbb{R}^{n\times n}$ denotes the graph structure. In this paper, the graph is undirected and unweighted. $\boldsymbol{Y}\in\mathbb{R}^{n\times 1}$ includes node labels from a class set $\mathcal{C}$.
@@ 3.2 Graph Neural Networks
Graph Neural Networks (GNNs) are tools for representation learning in node classification problems. The node representation in GNN is calculated by aggregating messages from neighbouring nodes. GNN can be represented as a function:
@@
where $\theta$ is the model parameter and $\boldsymbol{E}\in\mathbb{R}^{n\times d^{\prime}}$ denotes $d^{\prime}$-dimensional node embeddings. For example, the following is the output $\boldsymbol{E}^{(0)}$ of the fist layer of a typical Graph Convolutional Network (GCN):
@@
where $\boldsymbol{L}=\boldsymbol{D}^{-\frac{1}{2}}\boldsymbol{A}\boldsymbol{D}^{-\frac{1}{2}}$ is the normalised Laplacian adjacent matrix, and $\boldsymbol{D}$ is the degree matrix. $\boldsymbol{W}^{(0)}\in\mathbb{R}^{d\times d^{\prime}}$ is weights of the first convolutional layer. $\sigma(\cdot)$ is the activation operation for improving the model’s non-linearity.
@@ 3.3 Graph Condensation
Graph condensation aims to synthesis a small graph $\mathcal{\tilde{G}}=\{\boldsymbol{\tilde{A}},\boldsymbol{\tilde{X}},\boldsymbol{\tilde{Y}}\}$ for a large graph $\mathcal{G}=\{\boldsymbol{A},\boldsymbol{X},\boldsymbol{Y}\}$. The model trained with the synthetic graph is expected to have a similar performance as with the original graph. This objective is:
@@ 3.4 Node Classification in Continual Graph Learning
For solving the original node classification problem, GNN uses a full-connected layer to infer the nodes label. For example, after getting the embedding $\boldsymbol{E}^{(0)}$ in Eq. 2, the predicted labels can be inferred by
@@
In node classification of CGL, a model is required to handle $K$ tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{K}\}$. For the $k_{th}$ task $\mathcal{T}_{k}$, an incoming graph $\mathcal{G}_{k}$ arrives and the model needs to be updated with $\mathcal{G}_{k}$ while be tested on all previous graphs and the incoming graph. Following [9, 8, 10], the setting is transductive learning and can be easily extended to inductive learning.
CGL problem has two different continual settings, task incremental learning (task-IL) and class incremental learning (class-IL). In task-IL, the model is only required to distinguish nodes in the same task. While in class-IL, the model is required to classify nodes from all tasks together. Class-IL is more challenging, and this paper focuses on this setting while also reporting the overall performance under task-IL.
@@ 4 Methodology
This section describes the details of the proposed PseUdo-label guided Memory bAnk (PUMA) framework. The framework details are shown in Fig. 2. PUMA first condenses the incoming graph considering unlabelled nodes with pseudo labels. Once PUMA is updated, model initialises all weights and trains with PUMA from scratch.
@@ 4.1 Fast Graph Condensation by Distribution Matching
Condensation-based memory bank stores condensed synthetic graphs to approximate the historical data distribution. In this section, we develop a efficient graph condensation with distribution matching, aiming to maintain a similar data distribution for the synthetic data as the original data. This approach serves as a replayed graph generation method.
@@
where $\text{Dist}(\cdot,\cdot)$ function calculates the distance between two graphs. Using distribution matching, the distance between two graphs is measured in the embedding space, where both graphs are encoded by the same encoder. CaT encodes the original graph with a specific dimensional GNN encoder and randomly initialises the encoder weights for better generalisation performance. PUMA disentangles feature propagation and feature transformation the graph encoding. The incoming graph only propagates once to avoid calculate the feature aggregation of the incoming graph repeatedly but retain the structural information of the incoming graph naturally. Aggregated features $\boldsymbol{F}$ of the original graph can be calculated as a pre-processing:
@@
where $\boldsymbol{L}$ is the normalised Laplacian adjacent matrix, and $\boldsymbol{X}$ is the nodes features of the original graph. Feature transformation can use a wide multilayer perceptron with randomly initialised weights $\theta$ ($\text{MLP}_{\theta}$):
@@
where ${\mathcal{\tilde{G}}}_{k}^{*}=\{\boldsymbol{\tilde{X}}_{k}^{*},\boldsymbol{\tilde{Y}}_{k}^{*}\}$ is the optimal replayed graph with distribution close to the distribution of the incoming graph. Maximum mean discrepancy (MMD) is used to empirically calculate the distribution distance between two graphs. The objective is to find an optimal $\mathcal{\tilde{G}}_{k}$ for MMD:
$\ell_{\text{MMD}}=\sum_{c\in\mathcal{C}_{k}}r_{c}\cdot||\text{Mean}(\boldsymbol{E}_{k,c})-\text{Mean}(\boldsymbol{\tilde{E}}_{k,c})||^{2},$ (9)
where $\mathcal{C}_{k}$ is the set of classes of nodes in $\mathcal{G}_{k}$, $\boldsymbol{E}_{c,k}$ and $\boldsymbol{\tilde{E}}_{c,k}$ are the embedding matrix of the incoming graph and condensed graph, respectively, where all nodes’ labels are $c_{k}$, and $r_{c,k}=\frac{|\boldsymbol{E}_{c,k}|}{|\boldsymbol{E}_{k}|}$ is the class ratio for class $c_{k}$. ||$\cdot$|| ⋅ | is the number of rows in a matrix. $\text{Mean}(\cdot)$ is the mean vector of the node embeddings.

### v2

# PUMA: Efficient Continual Graph Learning for Node Classification with Graph Condensation

@@ Abstract
When handling streaming graphs, existing graph representation learning models encounter a catastrophic forgetting problem, where previously learned knowledge of these models is easily overwritten when learning with newly incoming graphs. In response, Continual Graph Learning (CGL) emerges as a novel paradigm enabling graph representation learning from static to streaming graphs. Our prior work, Condense and Train (CaT) [26] is a replay-based CGL framework with a balanced continual learning procedure, which designs a small yet effective memory bank for replaying data by condensing incoming graphs. Although the CaT alleviates the catastrophic forgetting problem, there exist three issues: (1) The graph condensation algorithm derived in CaT only focuses on labelled nodes while neglecting abundant information carried by unlabelled nodes; (2) The continual training scheme of the CaT overemphasises on the previously learned knowledge, limiting the model capacity to learn from newly added memories; (3) Both the condensation process and replaying process of the CaT are time-consuming. In this paper, we propose a PsUdo-label guided Memory bAnk (PUMA) CGL framework, extending from the CaT to enhance its efficiency and effectiveness by overcoming the above-mentioned weaknesses and limits. To fully exploit the information in a graph, PUMA expands the coverage of nodes during graph condensation with both labelled and unlabelled nodes. Furthermore, a training-from-scratch strategy is proposed to upgrade the previous continual learning scheme for a balanced training between the historical and the new graphs. Besides, PUMA uses a one-time prorogation and wide graph encoders to accelerate the graph condensation and the graph encoding process in the training stage to improve the efficiency of the whole framework. Extensive experiments on four datasets for the node classification task demonstrate the state-of-the-art performance and efficiency over existing methods. The code has been released in https://github.com/superallen13/PUMA.
@@ 1 Introduction
Graph representation learning aims to model the graph-structured data, typically in recommender systems [9], traffic prediction [49] and protein function prediction [55]. Generally, graphs are treated as static data in traditional graph representation learning, where a model is fixed once it has been trained. However, for many scenarios in the real world (e.g., social networks, citation networks and knowledge graphs), graph data are changing and evolving in a streaming manner [51, 5, 33]. Recently, continual graph learning (CGL) has emerged to handle the streaming graph by adapting the static graph neural networks (GNNs). In CGL, the most significant challenge is how to address a catastrophic forgetting problem, where a model easily forgets the previously learned knowledge while overemphasising on the incoming graphs [65, 23, 57, 26].
@@
Although our previous CaT framework [26] has demonstrated that the incoming graphs can be successfully condensed to an extremely small scale, for example, 0.1% of the original size, for the continual learning, it remains a few challenges in further improving the efficacy and efficiency. (1) Neglected unlabelled nodes: In the previous condensation process, nodes without label will be considered as the target for condensation rather than just supporting the message passing for the labelled nodes, potentially overlooking the valuable information from unlabelled nodes. (2) Imbalanced learned knowledge: Since previously generated memories will be repeatedly replayed during the continual learning, historical knowledge is easier to retain while the perception of new information in incoming graphs is hindered. Figure 1 shows this situation for CGL that the CaT model converges faster and better for the knowledge from previous Task 0 with a smaller loss than the newly incoming Task 1. This is due to the fact that neural networks are prone to learn shortcuts from training data [7]. (3) Low efficiency: Both the condensation and the CGL training are inefficient, because of repetitive computations, slow convergence speed of condensation and unnecessary message passing in the GNN training. The condensation process requires to repeatedly map the incoming graphs to different embedding spaces, which causes a large number of recalculations for message passing among node neighbourhoods. Besides, condensed graph converging is impeded by narrow graph encoders (e.g., encoders with 512-dimensional hidden layers). Furthermore, message aggregations of neighbours on the edge-free replayed memories are redundant, which hinders the efficiency of CGL model training.
@@
Figure 1: Loss and accuracy comparisons for the first two tasks when training the replayed graph of Task 1 while replaying the replayed graph of Task 0 at the first 50 steps after the replayed graph of Task 0 has been learned 500 steps. The continual (dotted lines) and retrained (solid lines) loss values for Task 0 (red) and Task 1 (blue) are shown, providing a visual representation of an optimisation problem.
@@
In light of the above challenges in the CaT method for CGL, this paper proposes a novel PseUdo-label guided Memory bAnk (PUMA) CGL framework to extend the CaT in a more effective and efficient manner. Firstly, the PUMA framework develops a pseudo-labelling to integrate data from unlabelled nodes, enhancing the informativeness of the memory bank, which addresses the issue of neglected unlabelled nodes. Secondly, to tackle the inflexible historical knowledge problem, a retraining strategy is devised. This involves initialising the entire model before replaying to balance the learned knowledge among different tasks for more effective decision boundaries as well as ensuring a more stable learning process. Figure 1 represents the optimisation process is smother, and the converge speed is quicker after using the retraining scheme. Lastly, the repetitive message passing calculation in the condensation process is simplified via a one-time propagation method, which aggregates messages from neighbouring nodes across the entire incoming graph one time, and it can be stored for reuse, significantly reducing the computational effort. Furthermore, wide graph encoders including more neurons are developed to accelerate convergence during condensation, enhancing the update efficiency of the condensed memories. Finally, to improve the training efficiency of the CGL model, the training of a multilayer perceptron (MLP) in lieu of message-passing GNNs is proposed. Due to its edge-free nature, PUMA enables the MLP to learn feature extractions and utilises a GNN to infer on the graphs with edges. These solutions collectively enhance the efficacy and efficiency of the graph condensation-based CGL framework.
@@ 2.1 Graph Neural Networks
Graph Neural Networks (GNNs) are effective tools for graph tasks [50, 20, 38, 32]. Recent GNNs mainly employ message passing paradigm for feature representation learning [16, 40, 48, 10]. More advanced GNN architectures, such as PC-Conv [19] and heterogeneous graphs [45, 44] are introduced for versatile representation learning.
@@ 2.2 Continual Graph Learning
Continual graph learning is a task for handling streaming graph data in different areas [17, 1, 28, 21, 54]. In graph area, CGL methods can be categorised into three branches: regularisation [23], replay-based [65, 57], and architecture-based [58] methods. TWP [23] preserves the topological information of historical graphs by adding regularisation. HPNs [58] redesigns the GNN architecture to 3-layer prototypes for representation learning. ER-GNN [65] integrates memory-replay by storing representative nodes. Sparsified subgraph memory (SSM) [57] stores the sparsified subgraphs by dropping edges and nodes randomly. SEM-curvature [59] employs balanced Forman curvature to maximise structural information within the subgraph. Structural shift risk mitigation (SSRM) [37] uses regularisation-based techniques to mitigate this issue. [39, 65] ensembles the graph replay and regularisation techniques to improve the model performance. Some recent work focuses on constructing advanced data structures, such as dynamic graphs [53] and temporal graphs [24, 22], to solve the incremental graph problem. Compared with the existing replay-based methods [54], which can hardly guarantee the quality of the memory bank when the size is limited, PUMA and CaT, generate a small yet informative memory bank in a differential manner, allowing the model to learn solely from these balanced replayed graphs, thus mitigating the issue of training imbalance.
@@ 2.3 Graph Condensation
Dataset condensation generates a small and synthetic dataset to replace the original dataset and to train a model with similar performance. Dataset condensation has been applied in computer vision [43, 62, 60, 41, 61, 2]. Recently, gradient matching has been applied to graph condensation, such as GCond [15], DosCond [14] and MCond [6]. DM [61] aims to learn synthetic samples which have similar distribution with the original dataset to mimic sampling methods [3, 46, 36]. GCDM [25] uses the distribution matching for graph condensation. SFGC [63] uses a training trajectory meta-matching scheme to condense the original graph into edge-free graphs. SGDD [52] broadcasts the structure information of the original graph to a condensed graph in the spectral graph theory perspective. GCondenser[27] uniformly benchmarks current GC methods. Recent attempts on computer vision have directly applied dataset condensation to continual learning [29, 47, 34, 35, 8]. CaT [26] applies graph condensation to graph continual learning while ensuring the balanced training.
@@ 3.1 Graph Neural Networks
For a node classification problem, a graph is denoted as $\mathcal{G}=\{\boldsymbol{A},\boldsymbol{X},\boldsymbol{Y}\}$, where $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ is the $d$-dimensional feature matrix for $n$ nodes, and the adjacency matrix $\boldsymbol{A}\in\mathbb{R}^{n\times n}$ denotes the graph structure. The graph is undirected and unweighted. $\boldsymbol{Y}\in\mathbb{R}^{n\times 1}$ are labels from a class set $\mathcal{C}$.
Graph Neural Networks (GNNs) are tools for representation learning. The node representation in GNN is calculated by message passing from neighbouring nodes as:
@@
where $\theta$ is the model parameter and $\boldsymbol{E}\in\mathbb{R}^{n\times d^{\prime}}$ are $d^{\prime}$-dimensional node embeddings. The output $\boldsymbol{E}^{(0)}$ of the first layer of a Graph Convolutional Network (GCN) is:
@@
where $\boldsymbol{L}=\boldsymbol{D}^{-\frac{1}{2}}\boldsymbol{A}\boldsymbol{D}^{-\frac{1}{2}}$ is the normalised Laplacian adjacent matrix, and $\boldsymbol{D}$ is the degree matrix. $\boldsymbol{W}^{(0)}\in\mathbb{R}^{d\times d^{\prime}}$ is the weight of the first GCN layer. $\sigma(\cdot)$ is the activation function.
@@ 3.2 Graph Condensation
Graph condensation aims to synthesis a small graph $\mathcal{\tilde{G}}=\{\boldsymbol{\tilde{A}},\boldsymbol{\tilde{X}},\boldsymbol{\tilde{Y}}\}$ from a large graph $\mathcal{G}=\{\boldsymbol{A},\boldsymbol{X},\boldsymbol{Y}\}$. The model trained with the synthetic graph is expected to have a similar performance as with the original graph:
@@ 3.3 Node Classification in Continual Graph Learning
To solve the original node classification problem, GNN uses a linear layer to infer the node label. For example, with the embedding $\boldsymbol{E}^{(0)}$ in Eq. 2, the labels can be predicted as:
@@
For node classification of CGL, a model is required to handle $K$ tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{K}\}$. For the $k_{th}$ task $\mathcal{T}_{k}$, an incoming graph $\mathcal{G}_{k}$ arrives and the model needs to update with $\mathcal{G}_{k}$ while testing on previous and the incoming tasks.
CGL problem has two different settings, task incremental learning (task-IL) and class incremental learning (class-IL). In task-IL, the model is only required to distinguish nodes in the same task. While in class-IL, the model is required to classify nodes from all tasks together. Class-IL is more challenging, and this paper focuses on this setting while also reporting the overall performance under task-IL.
@@ 4 Methodology
This section describes the details of the proposed PseUdo-label guided Memory bAnk (PUMA) framework. The framework details are shown in Fig. 2.
@@ 4.1 Fast Graph Condensation by Distribution Matching
Condensation-based memory bank stores condensed synthetic graphs to approximate the historical data distribution. In this section, we develop an efficient graph condensation with distribution matching, aiming to maintain a similar data distribution as the original data.
@@
where $\text{Dist}(\cdot,\cdot)$ calculates the distance between two graphs. Using distribution matching, the distance is measured in the embedding space, where both graphs are encoded by the same encoder. CaT encodes the original graph with a specific dimensional GNN encoder and randomly initialises the encoder weights for generalisation. PUMA disentangles feature propagation and feature transformation. The incoming graph only propagates once to avoid the feature aggregation of the incoming graph repeatedly but retaining the structural information. Aggregated features $\boldsymbol{F}$ of the original graph can be calculated as pre-processing:
@@
where $\boldsymbol{L}$ is the normalised Laplacian adjacent matrix, and $\boldsymbol{X}$ is the node feature of the original graph. Feature transformation is from a linear layer with random weights $\theta$ ($\text{MLP}_{\theta}$):
@@
where ${\mathcal{\tilde{G}}}_{k}^{*}=\{\boldsymbol{\tilde{X}}_{k}^{*},\boldsymbol{\tilde{Y}}_{k}^{*}\}$ is the optimal replayed graph with distribution close to the incoming graph. Maximum mean discrepancy (MMD) is used to empirically calculate the distribution distance between two graphs. The objective is:
$\ell_{\text{MMD}}=\sum_{c\in\mathcal{C}_{k}}\frac{|\boldsymbol{E}_{c,k}|}{|\boldsymbol{E}_{k}|}\cdot||\text{Mean}(\boldsymbol{E}_{k,c})-\text{Mean}(\boldsymbol{\tilde{E}}_{k,c})||^{2},$ (9)
where $\mathcal{C}_{k}$ is the set of classes in $\mathcal{G}_{k}$, $\boldsymbol{E}_{c,k}$ and $\boldsymbol{\tilde{E}}_{c,k}$ are the embedding matrices of the incoming graph and condensed graph, where all nodes’ labels are $c_{k}$. ||$\cdot$|| ⋅ | is the number of rows in a matrix. $\text{Mean}(\cdot)$ is the mean embeddings.